A screwdriver 9in.long is used to open a can of paint. The tipof the screwdriver is placed under the lip of the lid with the fulcrum 0.15in. from the lip. A force of 30lb is applied to the other end of the screwdriver.Find the force on the lip of the lid
Answer by mananth(16946)

A screwdriver 9in.long is used to open a can of paint. The tipof the screwdriver is placed under the lip of the lid with the fulcrum 0.15in. from the lip. A force of 30lb is applied to the other end of the screwdriver.Find the force on the lip of the lid
The formula for a lever is
load * distance on one side = load * distance on the other side of the fulcrum
Let the load on the lid be x
x*0.15 = 30* 8.85
0.15x=265.5
x= 265.5/0.15
x= 1770 lbs.